# Detecting and recognizing characters in Greek papyri with YOLOv8, DeiT and SimCLR

### 23 Jan 2024 v1

@@ Abstract
The capacity to isolate and recognize individual characters from facsimile images of papyrus manuscripts yields rich opportunities for digital analysis. For this reason the ‘ICDAR 2023 Competition on Detection and Recognition of Greek Letters on Papyri’ was held as part of the 17th International Conference on Document Analysis and Recognition. This paper discusses our submission to the competition. We used an ensemble of YOLOv8 models to detect and classify individual characters and employed two different approaches for refining the character predictions, including a transformer based DeiT approach and a ResNet-50 model trained on a large corpus of unlabelled data using SimCLR, a self-supervised learning method. Our submission won the recognition challenge with a mAP of 42.2%, and was runner-up in the detection challenge with a mean average precision (mAP) of 51.4%. At the more relaxed intersection over union threshold of 0.5, we achieved the highest mean average precision and mean average recall results for both detection and classification. We ran our prediction pipeline on more than 4,500 images from the Oxyrhynchus Papyri to illustrate the utility of our approach, and we release the results publicly in multiple formats.
@@
Character Detection, Character Recognition, Greek, Papyri, Oxyrhynchus
MSC:
68T10, 62M45
[…]
[inst1]organization=Melbourne Data Analytics Platform, University of Melbourne,addressline=Grattan St, city=Parkville, postcode=3010, state=VIC, country=Australia
[…]
[inst2]organization=Melbourne Centre for Data Science, University of Melbourne,addressline=Grattan St, city=Parkville, postcode=3010, state=VIC, country=Australia
{highlights}
First place in character recognition challenge
Second place in character detection challenge
Best recall and precision results for detection and recognition results for IoU 0.5
Releasing prediction results in multiple formats for 4500+ Oxyrhynchus Papyri images
@@ 1 Introduction
The challenge of publishing editions of manuscripts from the ancient world is immense. There are countless documents in libraries, museums and monasteries that have yet to be edited or published in a digital form. For example, the Oxyrhynchus Papyri, discovered by Bernard P. Grenfell and Arthur S. Hunt, contains approximately half a million papyri and papyrus fragments (Parsons, 2007) only a small fraction of which have been published.
Machine learning can assist in the tasks of editing and analyzing these collections. The ability to detect and recognize individual characters is especially useful and can be used for tasks such as assisting with approximate dating and identifying disiecta membra written by the same scribe Faigenbaum-Golovin et al. (2022). Automating the transcription of these documents also allows large corpora of scanned documents to be quickly searched for keywords or phrases of interest Alabau et al. (2014).
To compare and evaluate methods for this task, Mathias Seuret and colleagues ran the ‘ICDAR 2023 Competition on Detection and Recognition of Greek Letters on Papyri’ as part of the 17th International Conference on Document Analysis and Recognition (Seuret et al., 2023). The competition evaluated two tasks, one for detecting the bounding box for individual characters and the other for correctly recognizing the letter within each bounding box. This paper outlines our submission and the results of applying our method to published images of the Oxyrhynchus Papyri.
@@ 2.1 Competition Dataset
The training dataset provided for the competition consisted of 153 images from 108 Greek papyrus manuscripts which preserve text from the Iliad of Homer. Bounding boxes were added and annotated with the Greek letter for each character present. Apostrophes and periods were also annotated but ignored in evaluating the results. We divided the training dataset into five cross-validation partitions. Multiple images present in the same source manuscript were included in the same partition. The test dataset included 34 images from 31 manuscripts. Each submission required a JSON file in the COCO format to be uploaded to the competition site on CodaLab (Pavao et al., 2022).
@@ 2.2 Supplemental Data
To supplement the images in the training dataset, we included 4,533 images from volumes XV–LXXXII of the Oxyrhynchus Papyri (excluding volume XXVII). Images from these volumes included literary texts like in the competition dataset but it also included documentary texts. To help train models specifically for recognition, we used the second version of the AL-PUB dataset which contains 205,797 cropped images of individual characters taken from images in the Oxyrhynchus Papyri collection as part of the ‘Ancient Lives Project’ (Swindall et al., ). These images were labeled with the Greek letter from each image.
@@ 3.1 Detection and preliminary recognition
For character detection and preliminary recognition, we used YOLO (You Only Look Once) (Redmon et al., 2016). The YOLO neural network model uses a series of convolutional layers to predict bounding boxes and probabilities for each categorical class simultaneously. Initial detections are made on coarse cells produced from the convolutions and, if the same object is predicted in multiple cells, they are corrected using non-maximal suppression. The release of YOLO9000 (Redmon and Farhadi, 2017) and YOLOv3 (Redmon and Farhadi, 2018) brought about multiple enhancements. In 2020, Glen Jocher released a PyTorch implementation of YOLOv3 which he named YOLOv5, published by Ultralytics (Jocher, 2020). Our models used Jocher’s later version YOLOv8(Jocher et al., 2023). We used the ‘x’ sized pretrained model which is reported to have achieved a mAPval of 68.2 on the COCO val2017 dataset. We trained the model using the test dataset at three image resolutions: small (1280$\times$1280), medium (1600$\times$1600) and large (2048$\times$2048). Each YOLOv8 model was trained for 200 epochs and the weights from the epoch with the best result on the validation set were saved. The trained model was used to make predictions on the unannotated Oxyrhynchus Papyri images which were used as pseudo-labels (Lee et al., 2013) and concatenated with the competition training set. The YOLOv8 models were then trained using the combined input from the competition dataset and the pseudo-labels.
@@ 3.2 Recognition specific models
In the first instance, we trained a self-supervised SimCLR model (Chen et al., 2020a) on all of the available data containing cropped Greek characters. This included the competition dataset, the AL-PUB dataset, and the characters detected by the YOLOv8 models from the Oxyrhynchus Papyri.
The SimCLR approach takes a mini-batch of images as usual, but augments each image to produce two views and employs a contrastive loss to teach the network to match each augmented view in the batch with its positive pair. On datasets such as ImageNet and CIFAR-10, augmentations including cropping, horizontal flipping, color-jitter and Gaussian blur are used to get best results (Chen et al., 2020a).
Selecting appropriate augmentations for the problem is important for models trained using SimCLR to perform well because these influence the image features leveraged by the model to distinguish different images within the training process (Chen et al., 2021). In the context of Greek character detection, augmentations such as cropping and horizontal flipping may change the character sufficiently to make it a different class. Accordingly we reduced the degree of cropping, removed the horizontal flips and did stronger color-jitter and blur.
After pretraining a ResNet-50 convolutional neural network on the data, we fine-tuned the model (Chen et al., 2020b) using the labels available in the competition and AL-PUB datasets using fivefold cross-validation.
In the second instance, we trained DeiT transformer models using transfer learning (Touvron et al., 2021). This was an entirely supervised process so we used the same labeled data as for the ResNet-50 SimCLR models with the same cross-validation splits. For this approach we also used less aggressive cropping during training, so that a significant portion of the character of interest would always be present during the training process.
@@ 3.3 Ensembling
For our final submission, we ensembled the results from our trained models. For character detection, we ensembled the results for the YOLOv8 trained models using Weighted Boxes Fusion (Solovyev et al., 2021). We then cropped the characters resulting from this procedure and used our recognition models to make predictions for each. We first ensembled the predictions for the ResNet-50 SimCLR and DeiT-small models separately, using a hard majority voting approach. We then ensembled the YOLOv8 classification with the ResNet-50 SimCLR and DeiT-small overall votes, by adjusting the bounding box confidence by the proportion of votes for each method for a particular character.
@@ 4 Evaluation
The competition used the mean average precision (mAP) evaluation metric as used by the COCO dataset which averages results for ten intersection over union (IoU) values, with increments of 0.05 from 0.5 to 0.95. This was the primary metric which determined the winner for both the detection and recognition tasks. Results reported below use this metric unless stated otherwise. Also reported are the mean average precision and mean average recall (mAR) scores for an IoU value of 0.5, the traditional IoU value for evaluating object detection used in the pascal Visual Object Classes challenge (Everingham et al., 2010), and an IoU of 0.75 which requires a stricter localization of the bounding box.
@@ 7 Discussion
Since 1898, only a small fraction of the Oxyrhynchus Papyri has been catalogued and transcribed. This work, along with the character recognition models presented by Swindall et al. (2021) and the other submissions to the ICDAR competition Seuret et al. (2023), demonstrate the potential of new machine learning and AI approaches to make large scale analysis of these documents feasible.
The ability to reliably detect bounding boxes for characters in Greek manuscripts, and to recognize the letters within them, also opens many avenues for quantitative analysis. For example, similarities between the characters from different fragments might be used to discover disiecta membra of the same manuscript (such as P.Oxy XVIII.2170 and PSI XI.1218 (Lobel et al., 1941, 54)) or to identify separate manuscripts written by the same scribe (for a list of such instances, see in Johnson (2004), pp. 61–65).
Similarly, the isolated characters could be used to automatically detect the style of the hand. The isolated characters could be used as input to a machine learning model to predict the approximate date of the manuscript (for a survey of automated techniques for manuscript dating see Omayio et al. (2022)). The localization of the characters on the page can automate the process of identifying the widths and heights of the columns, similar to the analysis of Johnson (2004), pp. 162–212. The widths and heights of individual characters could also be used to refine probabilistic models for assessing the content that could have been in manuscript lacunae, such as that of McCollum (2022).
8 Conclusion
@@
9 Availability
10 Conflict of Interest
None.
11 Acknowledgements
This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. This project was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. We thank David Turnbull, Kamalpreet Singh and Daniel Russo-Batterham for assistance with different aspects of this project. Evelyn Mannix’s contribution to this research was supported by an Australian Government Research Training Program (RTP) Scholarship.
Supplementary Material
The confusion matrix from the cross validation results of the recognition specific models discussed in §5.1 is given in fig. 5.
Figure 5: Confusion matrix for character predictions on the AL-PUB and competition datasets, ensembling the SimCLR Resnet-50 and DeiT models. Results shown are the predictions obtained using the averaged probability from both methods on validation sets obtained using fivefold cross-validation.

### 14 Feb 2024 v2

@@ Abstract
Purpose: The capacity to isolate and recognize individual characters from facsimile images of papyrus manuscripts yields rich opportunities for digital analysis. For this reason the ‘ICDAR 2023 Competition on Detection and Recognition of Greek Letters on Papyri’ was held as part of the 17th International Conference on Document Analysis and Recognition. This paper discusses our submission to the competition.
Methods: We used an ensemble of YOLOv8 models to detect and classify individual characters and employed two different approaches for refining the character predictions, including a transformer based DeiT approach and a ResNet-50 model trained on a large corpus of unlabelled data using SimCLR, a self-supervised learning method.
Results: Our submission won the recognition challenge with a mAP of 42.2%, and was runner-up in the detection challenge with a mean average precision (mAP) of 51.4%. At the more relaxed intersection over union threshold of 0.5, we achieved the highest mean average precision and mean average recall results for both detection and classification.
Conclusion: The results demonstrate the potential for these techniques for automated character recognition on historical manuscripts. We ran the prediction pipeline on more than 4,500 images from the Oxyrhynchus Papyri to illustrate the utility of our approach, and we release the results publicly in multiple formats.
@@
character, detection, recognition, greek, papyri, oxyrhynchus
pacs:
MSC Classification]68T10, 62M45
@@ 1 Introduction
The challenge of publishing editions of manuscripts from the ancient world is immense. There are countless documents in libraries, museums and monasteries that have yet to be edited or published in a digital form. For example, the Oxyrhynchus Papyri, discovered by Bernard P. Grenfell and Arthur S. Hunt, contains approximately half a million papyri and papyrus fragments [14] only a small fraction of which have been published.
Machine learning can assist in the tasks of editing and analyzing these collections. The ability to detect and recognize individual characters is especially useful and can be used for tasks such as assisting with approximate dating and identifying disiecta membra written by the same scribe [6]. Automating the transcription of these documents also allows large corpora of scanned documents to be quickly searched for keywords or phrases of interest [1].
To compare and evaluate methods for this task, Mathias Seuret and colleagues ran the ‘ICDAR 2023 Competition on Detection and Recognition of Greek Letters on Papyri’ as part of the 17th International Conference on Document Analysis and Recognition [19]. The competition evaluated two tasks, one for detecting the bounding box for individual characters and the other for correctly recognizing the letter within each bounding box. This paper outlines our submission and the results of applying our method to published images of the Oxyrhynchus Papyri.
@@ 2.1 Competition Dataset
The training dataset provided for the competition consisted of 153 images from 108 Greek papyrus manuscripts which preserve text from the Iliad of Homer. Bounding boxes were added and annotated with the Greek letter for each character present. Apostrophes and periods were also annotated but ignored in evaluating the results. We divided the training dataset into five cross-validation partitions. Multiple images present in the same source manuscript were included in the same partition. The test dataset included 34 images from 31 manuscripts. Each submission required a JSON file in the COCO format to be uploaded to the competition site on CodaLab [15].
@@ 2.2 Supplemental Data
To supplement the images in the training dataset, we included 4,533 images from volumes XV–LXXXII of the Oxyrhynchus Papyri (excluding volume XXVII). Images from these volumes included literary texts like in the competition dataset but it also included documentary texts. To help train models specifically for recognition, we used the second version of the AL-PUB dataset which contains 205,797 cropped images of individual characters taken from images in the Oxyrhynchus Papyri collection as part of the ‘Ancient Lives Project’ [21]. These images were labeled with the Greek letter from each image.
@@ 3.1 Detection and preliminary recognition
For character detection and preliminary recognition, we used YOLO (You Only Look Once) [18]. The YOLO neural network model uses a series of convolutional layers to predict bounding boxes and probabilities for each categorical class simultaneously. Initial detections are made on coarse cells produced from the convolutions and, if the same object is predicted in multiple cells, they are corrected using non-maximal suppression. The release of YOLO9000 [16] and YOLOv3 [17] brought about multiple enhancements. In 2020, Glen Jocher released a PyTorch implementation of YOLOv3 which he named YOLOv5, published by Ultralytics [7]. Our models used Jocher’s later version YOLOv8[8]. We used the ‘x’ sized pretrained model which is reported to have achieved a mAPval of 68.2 on the COCO val2017 dataset. We trained the model using the test dataset at three image resolutions: small (1280$\times$1280), medium (1600$\times$1600) and large (2048$\times$2048). Each YOLOv8 model was trained for 200 epochs and the weights from the epoch with the best result on the validation set were saved. The trained model was used to make predictions on the unannotated Oxyrhynchus Papyri images which were used as pseudo-labels [10] and concatenated with the competition training set. The YOLOv8 models were then trained using the combined input from the competition dataset and the pseudo-labels.
@@ 3.2 Recognition specific models
In the first instance, we trained a self-supervised SimCLR model [2] on all of the available data containing cropped Greek characters. This included the competition dataset, the AL-PUB dataset, and the characters detected by the YOLOv8 models from the Oxyrhynchus Papyri.
The SimCLR approach takes a mini-batch of images as usual, but augments each image to produce two views and employs a contrastive loss to teach the network to match each augmented view in the batch with its positive pair. On datasets such as ImageNet and CIFAR-10, augmentations including cropping, horizontal flipping, color-jitter and Gaussian blur are used to get best results [2].
Selecting appropriate augmentations for the problem is important for models trained using SimCLR to perform well because these influence the image features leveraged by the model to distinguish different images within the training process [4]. In the context of Greek character detection, augmentations such as cropping and horizontal flipping may change the character sufficiently to make it a different class. Accordingly we reduced the degree of cropping, removed the horizontal flips and did stronger color-jitter and blur.
After pretraining a ResNet-50 convolutional neural network on the data, we fine-tuned the model [3] using the labels available in the competition and AL-PUB datasets using fivefold cross-validation.
In the second instance, we trained DeiT transformer models using transfer learning [22]. This was an entirely supervised process so we used the same labeled data as for the ResNet-50 SimCLR models with the same cross-validation splits. For this approach we also used less aggressive cropping during training, so that a significant portion of the character of interest would always be present during the training process.
@@ 3.3 Ensembling
For our final submission, we ensembled the results from our trained models. For character detection, we ensembled the results for the YOLOv8 trained models using Weighted Boxes Fusion [20]. We then cropped the characters resulting from this procedure and used our recognition models to make predictions for each. We first ensembled the predictions for the ResNet-50 SimCLR and DeiT-small models separately, using a hard majority voting approach. We then ensembled the YOLOv8 classification with the ResNet-50 SimCLR and DeiT-small overall votes, by adjusting the bounding box confidence by the proportion of votes for each method for a particular character.
@@ 4 Evaluation
The competition used the mean average precision (mAP) evaluation metric as used by the COCO dataset which averages results for ten intersection over union (IoU) values, with increments of 0.05 from 0.5 to 0.95. This was the primary metric which determined the winner for both the detection and recognition tasks. Results reported below use this metric unless stated otherwise. Also reported are the mean average precision and mean average recall (mAR) scores for an IoU value of 0.5, the traditional IoU value for evaluating object detection used in the pascal Visual Object Classes challenge [5], and an IoU of 0.75 which requires a stricter localization of the bounding box.
@@ 7 Discussion
Since 1898, only a small fraction of the Oxyrhynchus Papyri has been catalogued and transcribed. This work, along with the character recognition models presented by Swindall et al. [21] and the other submissions to the ICDAR competition [19], demonstrate the potential of new machine learning and AI approaches to make large scale analysis of these documents feasible.
The ability to reliably detect bounding boxes for characters in Greek manuscripts, and to recognize the letters within them, also opens many avenues for quantitative analysis. For example, similarities between the characters from different fragments might be used to discover disiecta membra of the same manuscript (such as P.Oxy XVIII.2170 and PSI XI.1218 [11, 54]) or to identify separate manuscripts written by the same scribe (for a list of such instances, see in [9], pp. 61–65).
Similarly, the isolated characters could be used to automatically detect the style of the hand. The isolated characters could be used as input to a machine learning model to predict the approximate date of the manuscript (for a survey of automated techniques for manuscript dating see [13]). The localization of the characters on the page can automate the process of identifying the widths and heights of the columns, similar to the analysis of [9], pp. 162–212. The widths and heights of individual characters could also be used to refine probabilistic models for assessing the content that could have been in manuscript lacunae, such as that of [12].
@@
[…]
Supplementary information
The confusion matrix from the cross validation results of the recognition specific models discussed in §5.1 is found in the supplementary material.
[…]
Acknowledgements
We thank David Turnbull, Kamalpreet Singh and Daniel Russo-Batterham for assistance with different aspects of this project.
Statements and Declarations
This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. This project was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. Evelyn Mannix’s contribution to this research was supported by an Australian Government Research Training Program (RTP) Scholarship.
[…]
Financial interests The authors declare they have no financial interests.
[…]
Non-financial interests None.
[…]
Data availability The training and validation data for the competition is available at https://codalab.lisn.upsaclay.fr/competitions/12419. Oxyrhynchus Papyri images are available at https://oxyrhynchus.web.ox.ac.uk/images. The AL-PUB version 2 dataset is available at https://data.cs.mtsu.edu/al-pub/. Our detections of characters from the Oxyrhynchus Papyri is available at https://doi.org/10.6084/m9.figshare.25011140.v1.